OpenStudy (anonymous):

Find the solution by the addition method.Check answers. A) -x+y= -3 -2x-y=6

OpenStudy (anonymous):

1-Multiply the first equation by 2 . ====> -2x+2y=-6 2- Add the two equations . -2x +2y=-6 + -2x - y=6 ========= 0 + y=0 y=0 3- substitute in any equation instead of y with 0 and get the value of x

OpenStudy (anonymous):

-x+y= -3 time -2 = 2x -2y = 6 -2x -y =6 ------------ 0x - 3y = 12 y = 12/-3 y = -4 substitution y= -4 than find for y -x+y= -3 -x-4 = -3 -4 + 3 = x -1 = x y = -4 , x = -1 http://www.wolframalpha.com/input/?i=-x%2By%3D+-3++%3B+-2x-y%3D6